Two hikers are 11 miles apart and walking toward each other. They meet in 2 hours. Find the rate of each hiker if one hiker walks 1.3 mph faster than the other.
Answer by rajagopalan(174)

Let the Speed of the slow hiker =x
Then Speed of fast hiker =x+1.3
Distance covered by them in 1 hr=2x+1.3 km
Distance covered by them in 2 hrs=2(2x+1.3)=4x+2.6 km
We have 4x+2.6=11
Then 4x=11-2.6
gives 4x=8.4
resulting in x=8.4/4=2.1 km
Speed of slow hiker = 2.1 km ph
Speed of fast hiker=3.4 km ph
Verify:
Distance covered by slow hiker in 2 hrs =4.2
Distance covered by fast hiker in 2 hrs =6.8
Summing upto.......................................=11.0 km
Hence OK.